please help me solve this equation
solve for y. provide the answer as a fraction if it is not a whole number.
1/8(16y+8)-17=-1/4(8y-16)

Answer by lwsshak3(11628)

1/8(16y+8)-17=-1/4(8y-16)
**
1/8(16y+8)-17=-1/4(8y-16)
(16y+8)/8-17=-(8y-16)/4
simplify
(2y+1)-17=-(2y-4)
2y-16=-2y+4
4y=20
y=5